yoshh at stanford.edu
Sat Oct 19 16:09:52 PDT 2002
I also run LTO400 yesterday successfully.
> As per your request, please, find below details of today's successfull
> SiH4: 14
> PH3: 86
> O2: 115
> Temp: 400
> Pressure: 326
> thickness: around 2000A
> Ankur Jain wrote:
> > Hello everyone!
> > I have a request to make to labmembers who will be using tylanbpsg in
> > next few days. It would be nice if they could post to this list the
> > pressures they get in their runs.
> > Thanks!
> > Ankur.
> > On Fri, 18 Oct 2002, Ted Berg wrote:
> > > hello all, we checked out the system this morning and found the pump
> > > screen clogged again. it was changed and tested at process flows:
> > > o2-115, siH4-100 pressure at 297 which is normal. please advise if
> > > have any further problems. also be aware that thickness is at 25+
> > > microns clean is scheduled for 30 so probably next week.ted
> > >
More information about the tylanbpsg